OpenStudy (anonymous):

please evaluate the following and show full working thank you -1/3 + 2 -1 5/6

OpenStudy (anonymous):

- 1/3 + 2 - 1 5/6 (convert mixed number to improper fraction) -1/3 + 2 - 11/6 (now find common denominator which is 6) -2/6 + 12/6 - 11/6 = - 13/6 + 12/6 = -1/6 <== answer